# Performance Estimation in Binary Classification Using Calibrated Confidence

### v1

@@ Abstract
Model monitoring is a critical component of the machine learning lifecycle, safeguarding against undetected drops in the model’s performance after deployment. Traditionally, performance monitoring has required access to ground truth labels, which are not always readily available. This can result in unacceptable latency or render performance monitoring altogether impossible. Recently, methods designed to estimate the accuracy of classifier models without access to labels have shown promising results. However, there are various other metrics that might be more suitable for assessing model performance in many cases. Until now, none of these important metrics has received similar interest from the scientific community. In this work, we address this gap by presenting CBPE, a novel method that can estimate any binary classification metric defined using the confusion matrix. In particular, we choose four metrics from this large family: accuracy, precision, recall, and F1, to demonstrate our method. CBPE treats the elements of the confusion matrix as random variables and leverages calibrated confidence scores of the model to estimate their distributions. The desired metric is then also treated as a random variable, whose full probability distribution can be derived from the estimated confusion matrix. CBPE is shown to produce estimates that come with strong theoretical guarantees and valid confidence intervals.
@@ 1 Introduction
In recent years, the problem of unsupervised accuracy estimation [Kivimäki et al., 2025] has drawn increasing attention. The task is to estimate the accuracy of a given machine learning (ML) model under some potentially shifted data distribution when measuring the accuracy directly is either inconvenient or altogether impossible due to lack of access to ground truth (GT) labels. Such situations are surprisingly common in real-life use cases where querying the labels for in-production data might be prohibitively expensive, or the labels might become available only after unacceptable latency [Krawczyk et al., 2017]. However, most users would still like to remain informed about the predictive performance of the deployed model.
Several different strategies for solving the unsupervised accuracy estimation problem have been suggested, such as training an ensemble of models and comparing the predictions of the models in the ensemble [Baek et al., 2022, Chen et al., 2021a, Jiang et al., 2022]. Another approach has tried to leverage importance weighting [Lu et al., 2022] to adjust the known performance of a source distribution on some unknown target distribution [Chen et al., 2021b]. Approaches leveraging distributional distances have also been suggested [Deng et al., 2023, Lu et al., 2023]. In this work, we focus on methods that utilize confidence scores produced by the model to estimate its performance under some target distribution [Deng et al., 2023, Garg et al., 2022, Guillory et al., 2021, Hendrycks and Gimpel, 2017]. Some of these methods have variants, where the uncertainty of the model is expressed in terms of prediction entropy instead of model confidence [Garg et al., 2022, Guillory et al., 2021].
A notable shortcoming of most earlier approaches has been that they all focus solely on predicting the accuracy of a given model. However, accuracy can be an uninformative or misleading metric in many cases, such as when there is a severe class imbalance or when different types of errors incur different costs [Bekkar et al., 2013]. In this work, we offer a remedy by expanding the scope of unsupervised accuracy estimation into unsupervised performance estimation and introducing a family of Confidence-based performance estimators (CBPE). Our approach applies to any classification metric, which can be defined using the elements of the Confusion Matrix. Since there is a plethora of such metrics, we introduce our method using only four common members of this family (accuracy, precision, recall, and F1) as examples.
@@
Finally, most earlier approaches have been examined only empirically in a limited number of settings and come with no theoretical guarantees. Alternatively, where such guarantees are given, they work only under strict conditions [Garg et al., 2022] or require extensive computational resources [Chen et al., 2021a]. In contrast, CBPE requires only that the estimated model’s confidence scores are calibrated. It has already been shown that under this calibration assumption CBPE yields unbiased and consistent estimates for model accuracy [Kivimäki et al., 2025]. In this paper, we extend these theoretical guarantees to other classification metrics. Furthermore, CBPE allows the user to derive the full probability distribution of the estimated metric, which can then be used to quantify uncertainty in the estimated metrics in the form of valid confidence intervals.
@@
    We present CBPE, a novel method for deriving estimates for different binary classification metrics from the estimated confusion matrix using calibrated confidence scores.
@@
    Each estimated metric comes in the form of a full probability distribution, and allows the user to quantify the uncertainty in the estimates in the form of valid confidence intervals.
@@
    For each metric, we show how to utilize a shortcut to derive or approximate the expected value of the distribution for situations where deriving the full distribution might be computationally too expensive. These shortcuts come with strong theoretical justifications.
@@ 2 Background
In this section, we present the necessary background information. We begin with calibration, then briefly discuss different types of distributional shifts, and finally, introduce the Poisson binomial distribution.
@@ 2.1 Confidence Calibration
Our estimators require that the model being monitored produces confidence scores and that these scores are or can be calibrated. Otherwise, the estimates might be biased. Let us now define precisely what we mean by calibration. Assume that the binary classification model $f$ under scrutiny generates a confidence score $S\in[0,1]$ in addition to its predictions $\hat{Y}=f(X)$, where the features $X$ and label $Y$ are sampled from a joint distribution $(X,Y)\sim p(\boldsymbol{x},y)$.
@@
Model $f$ is perfectly calibrated within distribution $p(\textbf{x},y)$ iff.
@@
In effect, this definition means that when a model is calibrated, any of its confidence scores can be interpreted as a probability that the predicted instance belongs to the positive class.
Since in practice, no model can achieve perfect calibration, it is useful to quantify how far from perfect the model is. How calibration error should be measured is an open research problem with many different approaches [Guo et al., 2017, Nixon et al., 2019, Posocco and Bonnefoy, 2021, Vaicenavicius et al., 2019]. In this paper, we use Adaptive Expected Calibration Error (ACE). In ACE, each prediction is assigned to one of $M$ bins based on its confidence score, where $M$ is a user-set hyperparameter. The bin boundaries are set adaptively so that each bin holds the same number of predictions, making ACE more robust regarding the choice of $M$ [Nixon et al., 2019]. ACE is defined as
$ACE=\sum_{m=1}^{M}\frac{|B_{m}|}{n}\left|A_{m}-C_{m}\right|,$ (1)
where $n$ is the total number of predictions, $B_{m}$ the set of predictions in bin $m$, and $A_{m}$ and $C_{m}$ the average accuracy and confidence of predictions in bin $m$ respectively. In this paper, we use $M=20$ bins in all experiments unless otherwise stated.
In our theoretical analyses of binary classifiers, we sometimes refer to confidence-consistency, which we will define next.
@@
Model $f$ is confidence-consistent in $p(\boldsymbol{x},y)$ iff. $\forall s\leavevmode\nobreak\ \in\leavevmode\nobreak\ [0,1]\leavevmode\nobreak\ \exists a\in\{0,1\}$: $P_{p(\boldsymbol{x},y)}(\hat{Y}=a\mid S=s)=1$.
@@ 2.2 Shifts in the Data Distribution
When training a supervised ML model $f$, we typically use data from a source distribution $p_{s}(\boldsymbol{x},y)$. However, once deployed, the model will encounter data from a potentially different target distribution $p_{t}(\boldsymbol{x},y)$, a scenario known as Dataset shift [Moreno-Torres et al., 2012]. If labels from the target distribution are inaccessible, the performance of $f$ under $p_{t}(\boldsymbol{x},y)$ can only be estimated.
Under these conditions, assumptions about the nature of the shift are necessary Garg et al. [2022]. Since $p(\boldsymbol{x},y)$ can be factorized as either $p(y|\boldsymbol{x})p(\boldsymbol{x})$ or $p(\boldsymbol{x}|y)p(y)$, the shift can be attributed to following components:
@@
2.3 Poisson Binomial Distribution
A fundamental building block in our considerations is the Poisson binomial distribution. It is a generalization of the binomial distribution, where the parameters $p_{i}$ for the $n$ Bernoulli trials are not necessarily equal. The probability mass function of a random variable $K$ following this distribution with parameters $p_{1},p_{2},...,p_{n}$ is
$P(K=k)=\sum_{A\in F_{k}}\prod_{i\in A}p_{i}\prod_{j\notin A}(1-p_{j}),$ (2)
where $F_{k}$ is the set of all subsets of $\{1,2,...,n\}$ with $k$ members. Since $|F_{k}|=\binom{n}{k}$, calculating the required probabilities using the definition directly is infeasible even for relatively small values of $n$. However, there is an efficient algorithm based on the Fast Fourier Transform with time complexity of only $\mathcal{O}(n\log n)$ [Hong, 2013]. In our experiments, we utilize an implementation of this algorithm found in the poibin [^1] https://github.com/tsakim/poibin python library.
@@ 3 Methodology
Throughout this section, we will assume that we are interested in monitoring a probabilistic binary classifier, which produces a prediction $\hat{Y}_{i}=f(X_{i})$ and a calibrated confidence score $S_{i}=s(X_{i})$ for each instance $X_{i}$. For simplicity, the monitored predictions are assumed to arrive in batches of fixed size $n$. Furthermore, we assume that we do not have access to actual labels $Y_{i}$ so we can not calculate any classification metrics directly, but have to rely on probabilistic estimates. We will denote the sets of indices for the positive and negative predictions as $I^{+}=\{i:\hat{Y}_{i}=1\}$ and $I^{-}=\{i:\hat{Y}_{i}=0\}$ respectively.
@@ 3.1 Estimating the Confusion Matrix
We begin by estimating the elements of the confusion matrix for a batch of $n$ predictions. Assume there are $n_{+}$ positive and $n_{-}$ negative predictions. Next, we will need to estimate how many of the positive and negative predictions actually belong to the positive and negative classes respectively.
Each positive prediction is either a true positive $(TP)$ or a false positive $(FP)$. Thus, it suffices to estimate the number of $TP$, since $|TP|+|FP|=n_{+}$. Assuming our model is calibrated, we can interpret a confidence score of a given prediction as the probability that the predicted instance belongs to the positive class. Thus, we consider the number of $TP$ to be a (discrete) random variable $X_{TP}$, which follows a Poisson binomial distribution, where the confidence scores for the positive predictions $\{S_{i}:i\in I^{+}\}$ act as parameters. If we are interested in having a point estimate for the number of $TP$, we can use the expected value of the distribution, $\mathbb{E}[X_{TP}]=\sum_{i\in I^{+}}S_{i}$, as such [^2] Note, that expectations such as this are conditional on the data $\{X_{i}\}_{i=1}^{n}$ and thus random variables until data is observed. However, we use $\mathbb{E}[X_{TP}]$ instead of more precise $\mathbb{E}[X_{TP}\mid\{X_{i}\}_{i=1}^{n}]$ for brevity.. The number of $FP$ is then a derived random variable $X_{FP}=n_{+}-X_{TP}$. Also, since
@@
Similarly, for negative predictions, the numbers of true negatives $(TN)$ and false negatives $(FN)$ can also be estimated using the confidence scores. As the calibrated confidence scores are probabilities that a given instance belongs to the positive class, the probability of a given instance $Y_{i}$ belonging to the negative class is the complement of its confidence score, formally, $P(Y_{i}=0)=1-S_{i}$. Thus, the number of $TN$ is a random variable $X_{TN}$, which again follows a Poisson binomial distribution, where the complements $1-S_{i}$ of the confidence scores $\{S_{i}:i\in I^{-}\}$ act as parameters. A point estimate for the number of $TN$ is then the expectation $\mathbb{E}[X_{TN}]=\sum_{i\in I^{-}}(1-S_{i})$. Finally, the number of $FN$ is a random variable $X_{FN}$ derived from $X_{FN}=n_{-}-X_{TN}$, with the expected value $\mathbb{E}[X_{FN}]=\sum_{i\in I^{-}}S_{i}$ The point estimates for the elements of the confusion matrix have been collected in Table 1.
@@
Let $f$ be a calibrated and confidence-consistent model operating on some target distribution $p_{t}(\boldsymbol{x},y)$. Then, $\bar{S}^{+}$, $\bar{S}^{-}$, $1-\bar{S}^{+}$, and $1-\bar{S}^{-}$ are unbiased and consistent estimators for $\operatorname{TPF}_{t}(f)$, $\operatorname{FPF}_{t}(f)$, $\operatorname{TNF}_{t}(f)$, and $\operatorname{FNF}_{t}(f)$ respectively.
@@ 3.2 Estimating Classification Metrics
In this paper, we devise algorithms for estimating four classification metrics used in binary classification: accuracy, precision, recall, and F1. These metrics can be defined with the elements of the confusion matrix (see Appendix A for specifics). We treat these metrics as random variables and derive their full distribution via their definitions. In what follows, we assume that we have already estimated the elements of the confusion matrix using the calibrated confidence scores from a given sample. After the full distribution of the derived random variable is known, one can use the expected value of the distribution as a point estimate for the metric of interest.
@@ 3.2.1 Accuracy
A naive way of forming the distribution of accuracy would be to first form the joint distribution of $X_{TP}$ and $X_{TN}$, which is straightforward since clearly $X_{TP}\perp\!\!\!\!\perp X_{TN}$, so the joint is just the product of the marginal distributions. After that, one could iterate over the (pairwise) values of the joint and collect the sums and their respective probabilities.
A simple shortcut is provided by realizing that
@@
where $X_{correct}$ can be derived in the following way: For all positive predictions, $S_{i}$ can directly be taken as the probability that the prediction was correct. On the other hand, for all negative predictions, one can interpret $1-S_{i}$ as the probability that the prediction was correct. Thus one can map the confidence scores $S_{i}\rightarrow Z_{i}$ with
@@
Then, $X_{correct}$ follows a Poisson binomial distribution with $Z_{i}$ as parameters. Taking the expected value of this distribution yields the Average Confidence (AC) [Hendrycks and Gimpel, 2017] estimator.

### v2

@@ Abstract
Model monitoring is a critical component of the machine learning lifecycle, safeguarding against undetected drops in the model’s performance after deployment. Traditionally, performance monitoring has required access to ground truth labels, which are not always readily available. This can result in unacceptable latency or render performance monitoring altogether impossible. Recently, methods designed to estimate the accuracy of classifier models without access to labels have shown promising results. However, there are various other metrics that might be more suitable for assessing model performance in many cases. Until now, none of these important metrics has received similar interest from the scientific community. In this work, we address this gap by presenting Confidence-based Performance estimation (CBPE), a novel method that can estimate any binary classification metric defined using the confusion matrix. In particular, we choose four metrics from this large family: accuracy, precision, recall, and F1, to demonstrate our method. CBPE treats the elements of the confusion matrix as random variables and leverages calibrated confidence scores of the model to estimate their distributions. The desired metric is then also treated as a random variable, whose full probability distribution can be derived from the estimated confusion matrix. CBPE is shown to produce estimates that come with strong theoretical guarantees and valid confidence intervals.
keywords:
artificial intelligence, machine learning, model monitoring, performance estimation
@@ 1 Introduction
In recent years, the problem of unsupervised accuracy estimation [kivimaki:2025] has drawn increasing attention. The task is to estimate the accuracy of a given machine learning (ML) model under some potentially shifted data distribution, when measuring the accuracy directly is either inconvenient or altogether impossible due to a lack of access to ground truth (GT) labels. Such situations are surprisingly common in real-life use cases where querying the labels for in-production data might be prohibitively expensive, or the labels might become available only after unacceptable latency [krawczyk:2017]. However, most users would still like to remain informed about the predictive performance of the deployed model.
A notable shortcoming of most earlier approaches has been that they all focus solely on predicting the accuracy of a given model. However, accuracy can be an uninformative or misleading metric in many cases, such as when there is a severe class imbalance or when different types of errors incur different costs [bekkar:2013]. In this work, we offer a remedy by expanding the scope of unsupervised accuracy estimation into unsupervised performance estimation and introducing Confidence-based performance estimation (CBPE). Our approach applies to any classification metric, which can be defined using the elements of the Confusion Matrix. Since there is a plethora of such metrics, we introduce our method using only four common members of this family (accuracy, precision, recall, and F1) as examples.
@@
Finally, most earlier approaches have been examined only empirically in a limited number of settings and come with no theoretical guarantees. Alternatively, where such guarantees are given, they work only under strict conditions [garg:2022] or require extensive computational resources [chen:2021]. In contrast, CBPE requires only that the estimated model’s confidence scores are calibrated. It has already been shown that under this calibration assumption CBPE yields unbiased and consistent estimates for model accuracy [kivimaki:2025]. In this paper, we extend these theoretical guarantees to other classification metrics. Furthermore, CBPE allows the user to derive the full probability distribution of the estimated metric, which can then be used to quantify uncertainty in the estimated metrics in the form of valid confidence intervals.
@@
    We present CBPE, a novel method for deriving estimates for different binary classification metrics using calibrated confidence scores.
@@
    Each estimated metric comes in the form of a full probability distribution, allowing the user to form valid confidence intervals for the estimates.
@@
    For each metric, we show a shortcut to derive or approximate the expected value in situations where deriving the full distribution might be computationally too expensive. These shortcuts come with strong theoretical justifications.
2 Related Works
Several different strategies for solving the unsupervised accuracy estimation problem have been suggested, such as training an ensemble of models and comparing the predictions of the models in the ensemble [baek:2022, chen:2021, jiang:2022]. Another approach has tried to leverage importance weighting [lu:2022] to adjust the known performance in a source distribution to some unknown target distribution [chen2021b]. Approaches leveraging distributional distances have also been suggested [deng:2023, lu:2023].
In this work, we focus on confidence-based estimators. These estimators utilize confidence scores produced by a model to estimate its performance under a given target distribution. Confidence scores are commonly used in ML models’ uncertainty estimation as a numerical indicator of a system’s certainty that its output is correct. Many types of models produce such scores innately [guo:2017, zadrozny:2001, zadrozny:2002], and there are ways to attach confidence scores to predictions of other models as well [kivimaki:2023]. Typically, the score is expressed on the interval $[0,1]$, but there are some exceptions to this [wei:2022].
Several confidence-based methods have been suggested, each exploiting different signals from model outputs. Average Confidence (AC) [hendrycks:2017, kivimaki:2025] provides a simple baseline by using the average of model’s confidence scores as an estimate for accuracy. Difference of Confidences (DoC) [guillory:2021] builds on this by training a simple regressor to map the difference of average confidences between source and target distributions as an estimate for model performance under the target distribution. Similarly, Average Thresholded Confidence (ATC) [garg:2022] learns thresholds that transfer across domains, allowing accuracy estimation on unseen distributions without labels. Finally, Nuclear Norm [deng:2023] analyzes both the confidence and diversity of prediction distributions across unlabeled data, aiming to better characterize error patterns. Some of these methods have variants, where the uncertainty of the model is expressed in terms of prediction entropy instead of model confidence [garg:2022, guillory:2021].
All of the above estimators can only be used to estimate model accuracy. As already briefly mentioned in Section 1, accuracy is a problematic metric in many cases. For example, with imbalanced data, high accuracy can be achieved trivially by always predicting the majority class. To the best of our knowledge, no confidence-based estimator that can estimate any other performance metric has been proposed prior to the method presented in this paper.
@@ 3 Background
In this section, we present the necessary background information. We begin with calibration, then briefly discuss different types of distributional shifts.
@@ 3.1 Confidence Calibration
Let us assume that the binary classification model $f$ under scrutiny generates a confidence score $S\in[0,1]$ in addition to its predictions $\hat{Y}=f(X)$, where the features $X$ and label $Y$ are sampled from a joint distribution $(X,Y)\sim p(\boldsymbol{x},y)$. Then, we have
@@
Model $f$ is perfectly calibrated within distribution $p(\boldsymbol{x},y)$ iff.
@@
In effect, this definition means that when a model is calibrated, any of its confidence scores can be interpreted as a probability that the predicted instance belongs to the positive class. In our theoretical analyses of binary classifiers, we sometimes refer to confidence-consistency, which we define as
@@
Model $f$ is confidence-consistent in $p(\boldsymbol{x},y)$ iff. $\forall s~\in~[0,1]~\exists a\in\{0,1\}$: $P_{p(\boldsymbol{x},y)}(\hat{Y}=a\mid S=s)=1$.
@@ 3.2 Shifts in the Data Distribution
When training a supervised ML model $f$, we typically use data from a source distribution $p_{s}(\boldsymbol{x},y)$. However, once deployed, the model will encounter data from a potentially different target distribution $p_{t}(\boldsymbol{x},y)$, a scenario known as Dataset shift [moreno-torres:2012]. If labels from the target distribution are inaccessible, the performance of $f$ under $p_{t}(\boldsymbol{x},y)$ can only be estimated. Under these conditions, assumptions about the nature of the shift are necessary [garg:2022]. Since $p(\boldsymbol{x},y)$ can be factorized as either $p(y|\boldsymbol{x})p(\boldsymbol{x})$ or $p(\boldsymbol{x}|y)p(y)$, the shift can be attributed to following components:
@@
Intuitively, concept shift means that the feature distribution remains unchanged, but the conditional distribution of labels changes. This is typical in adversarial settings, such as fraud detection, or in situations where model predictions affect user behavior, such as in recommender systems. In covariate shift, the opposite happens. This is typical in situations where the relation between features and labels rarely changes, such as when modelling diseases based on symptoms. Such a model might be trained on some (unrepresentative) subpopulation but applied to the whole population. Alternatively, during an epidemic, the prevalence of a disease in a fixed population might change over time, making it a case of label shift.
@@ 4 Methodology
Throughout this section, we are interested in monitoring a probabilistic binary classifier, which produces a prediction $\hat{Y}_{i}=f(X_{i})$ and a calibrated confidence score $S_{i}=s(X_{i})$ for each instance $X_{i}$. For simplicity, the monitored predictions are assumed to arrive in batches of fixed size $n$. Furthermore, we assume that we do not have access to actual labels $Y_{i}$ so we can not calculate any classification metrics directly, but have to rely on probabilistic estimates. We will denote the sets of indices for the positive and negative predictions as $I^{+}=\{i:\hat{Y}_{i}=1\}$ and $I^{-}=\{i:\hat{Y}_{i}=0\}$, respectively.
@@ 4.1 Estimating the Confusion Matrix
Our approach begins by estimating the elements of the confusion matrix for a batch of $n$ predictions. Assume there are $n_{+}$ positive and $n_{-}$ negative predictions. Next, we will need to estimate how many of the positive and negative predictions actually belong to the positive and negative classes, respectively.
Let us start with the positive predictions. Each positive prediction is either a true positive $(TP)$ or a false positive $(FP)$. Thus, it suffices to estimate the number of $TP$, since $|TP|+|FP|=n_{+}$. Since our model is assumed to be perfectly calibrated, we can interpret a confidence score of a given prediction as the probability that the predicted instance belongs to the positive class. Thus, we consider the number of $TP$ to be a (discrete) random variable $X_{TP}$, which follows a Poisson binomial distribution, where the confidence scores for the positive predictions $\{S_{i}:i\in I^{+}\}$ act as parameters. If we are interested in having a point estimate for the number of $TP$, we can use the expected value of the distribution, $\mathbb{E}[X_{TP}]=\sum_{i\in I^{+}}S_{i}$, as such [^1] Note, that expectations such as this are conditional on the data $\{X_{i}\}_{i=1}^{n}$ and thus random variables until data is observed. However, we use $\mathbb{E}[X_{TP}]$ instead of more precise $\mathbb{E}[X_{TP}\mid\{X_{i}\}_{i=1}^{n}]$ for brevity.. The number of $FP$ is then a derived random variable $X_{FP}=n_{+}-X_{TP}$. Also, since
@@
Similarly, for negative predictions, the numbers of true negatives $(TN)$ and false negatives $(FN)$ can also be estimated using the confidence scores. As the calibrated confidence scores are probabilities that a given instance belongs to the positive class, the probability of a given instance $Y_{i}$ belonging to the negative class is the complement of its confidence score, formally, $P(Y_{i}=0)=1-S_{i}$. Thus, the number of $TN$ is a random variable $X_{TN}$, which again follows a Poisson binomial distribution, where the complements $1-S_{i}$ of the confidence scores $\{S_{i}:i\in I^{-}\}$ act as parameters. A point estimate for the number of $TN$ is then the expectation $\mathbb{E}[X_{TN}]=\sum_{i\in I^{-}}(1-S_{i})$. Finally, the number of $FN$ is a random variable $X_{FN}$ derived from $X_{FN}=n_{-}-X_{TN}$, with the expected value $\mathbb{E}[X_{FN}]=\sum_{i\in I^{-}}S_{i}$ .
The point estimates for the elements of the confusion matrix have been collected in Table 1.
@@
Let $f$ be a perfectly calibrated and confidence-consistent model operating on some target distribution $p_{t}(\boldsymbol{x},y)$. Then, $\bar{S}^{+}$, $\bar{S}^{-}$, $1-\bar{S}^{+}$, and $1-\bar{S}^{-}$ are unbiased and consistent estimators for $\operatorname{TPF}_{t}(f)$, $\operatorname{FPF}_{t}(f)$, $\operatorname{TNF}_{t}(f)$, and $\operatorname{FNF}_{t}(f)$, respectively.
@@ 4.2 Estimating Classification Metrics
In this paper, we devise algorithms for estimating four classification metrics used in binary classification: accuracy, precision, recall, and F1. We treat these metrics as random variables and derive their full distribution. While the expressions for these derived random variables follow straightforwardly from the standard definitions, treating the estimated metrics as random variables allows us to capture the whole distribution of the estimated metric, which enables us to derive valid confidence intervals for the estimates. Furthermore, deriving the exact distributions using these expressions might require non-trivial calculations for some metrics. We will present algorithms for two such metrics, namely, recall and F1.
First, we need to estimate the elements of the confusion matrix using the calibrated confidence scores from a given sample. In what follows, we assume that this has already been done. Similarly to the estimated elements of the confusion matrix, after the full distribution of the derived random variable is known, one can use the expected value of the distribution as a point estimate for the metric of interest.
@@ 4.2.1 Accuracy
A naive way of forming the distribution of accuracy would be to first form the joint distribution of $X_{TP}$ and $X_{TN}$, which is straightforward since clearly $X_{TP}\perp\!\!\!\!\perp X_{TN}$, so the joint is just the product of the marginal distributions. After that, one can iterate over the (pairwise) values of the joint and collect the sums and their respective probabilities. A simple shortcut is provided by realizing that
@@
where $X_{correct}$ can be derived in the following way: For all positive predictions, $S_{i}$ can directly be taken as the probability that the prediction was correct. On the other hand, for all negative predictions, one can interpret $1-S_{i}$ as the probability that the prediction was correct. Thus, one can map the confidence scores $S_{i}\rightarrow Z_{i}$ with
@@
Then, $X_{correct}$ follows a Poisson binomial distribution with $Z_{i}$ as parameters. Taking the expected value of this distribution yields the Average Confidence (AC) [guillory:2021, hendrycks:2017] estimator, which has already been shown to be an unbiased and consistent estimator of model accuracy under the calibration assumption [kivimaki:2025].